### v1

# Dynamic and Adaptive Feature Generation with LLM

Abstract
The representation of feature space is a crucial environment where data points get vectorized and embedded for upcoming modeling. Thus the efficacy of machine learning (ML) algorithms is closely related to the quality of feature engineering. As one of the most important techniques, feature generation transforms raw data into an optimized feature space conducive to model training and further refines the space. Despite the advancements in automated feature engineering and feature generation, current methodologies often suffer from three fundamental issues: lack of explainability, limited applicability, and inflexible strategy. These shortcomings frequently hinder and limit the deployment of ML models across varied scenarios. Our research introduces a novel approach adopting large language models (LLMs) and feature-generating prompts to address these challenges. We propose a dynamic and adaptive feature generation method that enhances the interpretability of the feature generation process. Our approach broadens the applicability across various data types and tasks and draws advantages over strategic flexibility. A broad range of experiments showcases that our approach is significantly superior to existing methods.
00footnotetext: *Corresponding Authors.
1 Introduction
The success of machine learning (ML) algorithms generally depends on three aspects: data processing, feature engineering, and modeling. Among these, feature engineering is a critical part that directly influences the performance and effectiveness of ML models. Within feature engineering, feature generation is a vital process that transforms raw features into a structured format and optimizes the feature space by creating new features from original ones. The transformation usually involves mathematical or algorithmic operations on features. This optimization can significantly enrich the feature space, with which the ML models can draw and utilize the data information more effectively and achieve better model performance.
Given the importance of feature engineering and feature generation, there has been much work on enhancing both the effectiveness and efficiency of these methods. As deep machine learning explosively develops, automated feature engineering appears and becomes mainstream due to its convenience and remarkable performance. These auto-engineering algorithms greatly reduce the need for human calculation and determination and thus reduce subjective errors and mistakes. And in the field of automated feature generation there are outstanding works such as (Khurana et al., 2018) and (Wang et al., 2022). These methods leverage advanced techniques like reinforcement learning to generate optimal feature sets that enhance downstream tasks’ performance while maintaining a degree of explainability. Despite these methods enriching the feature space and restructuring it into a more refined space structure, like other automated feature generation and broader automated feature engineering techniques, these works have not resolved three fundamental issues:1) Lack of Explainability: Most feature engineering processes that utilize deep learning algorithms are result-oriented “black boxes”. The specific operations on features and the process of restructuring feature space are usually hard for researchers to learn and follow. 2) Limited Applicability: For different data types and downstream tasks, researchers need to manually select and switch among a wide range of feature engineering methods, which demands a high level of domain knowledge and expertise of the researchers. This limited applicability can be a significant barrier across varied ML scenarios. 3) Fixed Strategy: Existing methods usually require predefined strategies for exploration and generation. Models utilizing these methods are unable to adapt or refine their strategies based on insights acquired during the learning process of the feature space. Hence, these models lack the flexibility to dynamically adjust to the evolving understanding of the data.
Figure 1: Our goal is to iteratively reconstruct the feature space to create an optimal and explainable feature set that enhances performance in downstream machine learning tasks.
Our Target. We target on developing an automated feature generation method that: (1) has transparent and understandable feature generation process; (2) adapts to various data types and downstream tasks and selects best suiting operations automatically; (3) continuously learns from the feature space and refine its strategies based on new insights. For target (1) we clearly document the operations applied during the feature generation and how they transform the input data into a structured feature space. By doing so, we intend to replace the "black box" nature with a more open and interpretable framework. For target (2) we turn to large language model (LLM) that can assess the characteristics of the data and the requirements of the task and dynamically select the most appropriate strategies under proper guidance. For target (3) we develop multiple expert-level LLM agents with different strategies. These agents adjust their operations and strategies in response to changes in data patterns, task requirements, and actions of other agents.
Our Method. We employ a novel approach to automate and enhance the feature generation process by adopting LLM with expert-level guidance. Our method starts by inputting an original feature set and a predefined operation set into the LLM, which then creates several expert agents. Each agent generates new features with its exclusive strategy and integrates these new features with the original features to produce enriched feature sets. The newly created feature sets are applied to downstream tasks to evaluate their performance, and the performance outcomes are fed back to the agents. The agents “communicate” and share their inferences, and reflect on the impact and effectiveness of their own and other agents’ operations on the downstream tasks. In this way the agents refine and optimize their generation strategies to better capture how variations in the feature space impact downstream tasks, furthermore reconstruct the feature space more adaptively and effectively. The iteration continues with new feature sets until the optimal feature set for the task is found or a predefined number of iterations is reached.
Contributions. In summary, our main contributions include:
  • We propose LFG, a novel automated feature generation methodology based on large language model, to restructure feature space of a dataset. This end-to-end structure allows for training with LLM agents on varied feature generation strategies without manual selection of feature operations.
  • We explore a dynamic and adaptive generation process based on feedback from downstream tasks. This method effectively utilizes the in-context learning and reasoning capabilities of LLMs in the feature generation process and significantly enhances the applicability and effectiveness of the generated features across various machine learning scenarios without the need to create additional ML models.
  • We conduct a series of experiments to validate the effectiveness and robustness of our method (LFG) across different datasets and downstream tasks. Our experiments show that our method has clear advantages over existing methods and extensive potential of promoting to a broader range of feature engineering tasks.
2 Preliminary and Related Works
2.1 Feature Generation
Feature Engineering. Feature engineering is the process of selecting, modifying or creating new features from raw data to improve the performance of machine learning models Severyn and Moschitti (2013). The goal is to reconstruct an optimal and explainable feature representation space for a certain machine learning task (Zheng and Casari, 2018; Chen et al., 2021). This process can be described as a function ϕ::italic-$\phi:\mathcal{F}\rightarrow\mathcal{F}^{{}^{\prime}}$ that transforms the original feature set $\mathcal{F}$ into a new feature set $\mathcal{F}^{{}^{\prime}}$ with certain operations. Feature generation is a decisive category in the field of feature engineering. Feature generation means generating new attributes from existing features in a dataset through various mathematical or logical transformation operations. Feature generation methods can primarily be categorized into two main classes: (1) latent representation learning based methods, such as deep factorization machines methods (Guo et al., 2017; Song et al., 2019) and deep representation learning methods (Zhong et al., 2016; Bengio et al., 2013). These methods can create complex latent feature spaces; however, their generation processes often lack transparency and are difficult to trace and explain (Schölkopf et al., 2021). (2) feature transformation-based methods, which generate new features by performing arithmetic or aggregation operations (Nargesian et al., 2017). These methods often require a large amount of manual operations and highly rely on domain knowledge to select appropriate transformations and adjust parameters.
Automated Feature Generation. With the widespread adoption of large model techniques and various deep learning methods in feature engineering field, automated feature generation has significant development and application. Automated feature generation enhances the feature space by systematically creating and integrating new features that improve the model performance (Xiang et al., 2021; Wang et al., 2022; Pan et al., 2020; Wang et al., 2022). For example,  (Katz et al., 2016) developed ExploreKit, which generates a large set of candidate features by combining information in the original features.  (Khurana et al., 2016) explores the feature space using manually crafted heuristic traversal strategies, while  (Shi et al., 2018) proposed a feature optimization approach using deep learning and feature selection to enhance traffic classification performance. These methods are more efficient than traditional manual feature engineering and enable faster processing of large datasets. However, despite their efficiency and ease of operation, these automated feature generation methods often ignore the semantic aspects of data, and their “black box” operation process make the results even harder to explain.
2.2 Large Language Models (LLMs) and Tree of Thoughts
LLMs Capabilities. LLMs have revolutionized numerous fields with their extraordinary capabilities of in-context learning and step-by-step reasoning (Brown et al., 2020; Achiam et al., 2023; Wei et al., 2022a, b). With these capabilities, LLMs can effectively understand and navigate complex tasks with appropriate guidance and precise prompts. Thus they can exhibit outstanding professional competence across various specialized domains (Wang et al., 2023; White et al., 2023). Recent advancements in prompt engineering and reasoning structures have further enabled these models to tackle complex data analysis tasks and multi-step feature engineering processes (Peng et al., 2023; Liu et al., 2023).
LLM Prompting. LLM prompting leverages pre-train models to recall existing knowledge and generate output based on specific guideline prompts (White et al., 2023). LLM prompting interacts closely with LLMs’ capabilities of in-context learning and step-by-step reasoning (Giray, 2023; Chu et al., 2023). Prompts with context details guide the models in thinking more coherently when handling long tasks. Moreover, step-by-step reasoning prompts can guide the model through a logical sequence of thought steps by mimicking the human-like reasoning process of solving complicated problems (Yu et al., 2023; Wei et al., 2022b; Diao et al., 2023).
Tree of Thoughts. In prompting methods, one category is "logical reasoning prompting," which employs structured prompts to guide LLMs through logical analysis and problem-solving tasks. Input-Output (IO) Prompting (Wang et al., 2024; Hao et al., 2023) involves wrapping the input $x$ with task-specific instructions or examples to guide the model to produce the desired output $y$. This method is straightforward but its guiding capability is limited in complex tasks. Chain-of-Thought (CoT) Prompting (Wei et al., 2022b; Besta et al., 2024) involves a sequence of intermediate, coherent language expressions $z_{1},\ldots,z_{n}$ that logically bridge the input $x$ to the output $y$. This method enhances the model’s logical capabilities for complex problem-solving but it is limited to linear, single-strategy reasoning, which may only capture some possible solutions. Tree of Thoughts (ToT) Prompting (Yao et al., 2024) extends the CoT prompting by exploring multiple reasoning paths over thoughts. ToT works as a search over a tree structure where each node represents a partial solution within the input and contextual thoughts. ToT advantages over other methods by exploring multiple reasoning paths with different strategies. Thus it performs better in deep data analysis and multi-step feature engineering.
3 Methodology
In this section, we design a novel feature generation method called LLM Feature Generation (LFG), that is highly interpretable, dynamically adaptable, and employs an end-to-end approach to enhance the performance of downstream tasks significantly. In this method, we utilize a LLM as the auto-generator of features. We assist the LLM with the ToT technique to generate detailed interference along with each step of generation so that each decision is explainable. Furthermore, we collect the performance of downstream tasks to offer feedback to the LLM generator. Our entire workflow can be divided into two main parts: (1) Automated Feature Generation with LLM Agents; (2) Feedback and Optimization.
Figure 2: The framework of LFG. First, we input the original feature set and operation set into LLM windows. Second, we guide the LLM in creating three expert agents. Each agent generates new features with operations from the operation set, and then combines new features with original features to create a new feature subset. Then, we fed the updated feature subset into a downstream task. After that, we provide the performance of the feature subset in downstream tasks as feedback to agents and iterate the process until the best feature subset is found or the maximum number of iterations is reached.
3.1 Important Definitions
We first define the feature set and operation set along with their corresponding mathematical symbols.
Feature Set. Let $\mathcal{D}=\{\mathcal{F},\mathcal{y}\}$ be the dataset. Here, =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{n}\}$ is the feature set, where each column represents a feature $f_{i}$, and each row represents a data sample; $y$ is the corresponding target label set for the samples. Our main purpose is to reconstruct the feature space of the dataset $\mathcal{D}=\langle\mathcal{F},\mathcal{y}\rangle$.
Operation Set. The operation set $O$ consists of mathematical operations performed on existing features to generate new ones. Operations are of two types: unary and binary. Unary operations include “square”, “exp”, “log”, etc.; binary operations include “plus”, “multiply”, “divide”, etc. The mathematical definitions of specific operations are provided in the  Appendix A.2.
Optimization Objective. We aim to construct an optimal and interpretable feature space that can enhance the performance of downstream tasks. With the feature set $\mathcal{F}$ and the operation set $\mathcal{O}$, our optimization objective is to find a reconstructed feature set $\hat{\mathcal{F}}$:
$\mathcal{F}^{*}=\underset{\hat{\mathcal{F}}}{\text{argmax}}\;\mathbf{\theta}_{\mathcal{R}}(\hat{\mathcal{F}},\mathcal{y}),$ (1)
where $\mathcal{R}$ is a downstream ML task (e.g., classification, regression, ranking, detection), $\mathbf{\theta}$ is the performance indicator of $\mathcal{R}$ and $\hat{\mathcal{F}}$ is the update of feature set $\mathcal{F}$. Here, $\hat{\mathcal{F}}$ is generated by applying operations from $\mathcal{O}$ to the original feature set $\mathcal{F}$ using a certain algorithmic structure.
3.2 Automated Feature Generation with LLM Agents
In this section, we introduce LLM agents and provide details on feature generation. We utilize LLM to generate several agents for feature-generation tasks. Here, each agent acts as an automated feature-creator, which applies specific operations on a given feature set to generate new features or delete existing features according to the prompts given in each iteration. These agents can emulate expert-level logical thinking and decision-making capabilities, hence optimizing the feature set for input data of downstream tasks and enhancing tasks’ performance.
Agents. We first define an agent $\mathcal{A}_{l}$ which operates on a feature subset $\mathcal{F}_{l}=\{f_{1},\ldots,f_{n}\}$ by applying operations to the features $f_{i}$ and $f_{j}$. The $\mathcal{F}_{l}$ is the subset of $\mathcal{F}$. The agent draws operations from an operation set $\mathcal{O}$, and these operations generate new features:
$\mathcal{F}_{l}^{{}^{\prime}}=\{\mathcal{F}_{l},g_{1},\ldots,g_{k},\ldots\}.$ (2)
Each new feature $g_{k}$ is produced by applying an operation $o$ to the features $f_{i}$ and $f_{j}$:
$g_{k}=o(f_{i},f_{j})\quad\text{for each}\quad o\in\mathcal{O}.$ (3)
Here, $\mathcal{F}_{l}$ represents the initial feature subset given to $\mathcal{A}_{l}$, $\mathcal{F}_{l}^{{}^{\prime}}$ is the output feature subset encompassing the original and all new features generated by $\mathcal{A}_{l}$, and $g_{k}$ is the new feature generated by applying the operation $o$ to features $f_{i}$ and $f_{j}$. In this way we complete one generation node $s_{l}$ of $\mathcal{A}_{l}$, and a layer $Layer_{1}$ of nodes is of total $l$ generation nodes.
3.3 Feedback and Optimization
We now have a new feature set composed of all $l$ subsets, denoted as $\mathcal{F}_{l}^{{}^{\prime}}={\{\mathcal{F}_{l},g_{1},\ldots,g_{k}\}}$. After the agents generate all $l$ subsets, we move forward to evaluate the performance of these subsets in downstream tasks and the effectiveness of the generated nodes at this layer. We then gather the feedback of downstream tasks by the performance metrics $\mathbf{\theta}=\{\theta_{t}\}$, for each iteration $t$ corresponding to each generation layer. This feedback reflects the effectiveness of the newly created feature subset’s effectiveness and illustrates the agents’ contribution to the task. The agents evaluate their performance by comparing the metric $\theta_{t}$ with $\theta_{(t-1)}$. If $\theta_{t}$ significantly improves over $\theta_{(t-1)}$, the feature generation strategy is considered “effective”. The agents’ self-evaluation positively correlates with the increase in these metrics; a larger improvement signifies a more successful feature generation strategy.
Furthermore, each agent shares their thought process and inferences when generating features, including the logic behind selecting specific operations. This transparency helps agents understand and adopt diverse strategies to bring collaborative improvement. Through self-evaluation and peer interactions, agents identify which strategies successfully enhanced performance and which may require adjustments or replacements. In the next round, agents generate increasingly effective feature subsets adopting $\mathcal{F}_{l}^{{}^{\prime}}$ and $\mathcal{O}$ to form the next layer of generated nodes. The iterative process continues until the optimal feature set $\hat{\mathcal{F}}$ is found or the maximum number of iterations $T$ is reached:
^$\hat{\mathcal{F}}=\bigcup_{l\in L}\mathcal{F}_{l}^{\text{opt}},$ (4)
where $\mathcal{F}_{l}^{\text{opt}}$ is the optimal subset within each group $l$, and $L$ is the set of all groups or iterations.
Monte Carlo Tree Search. In order to further optimize the feature space and balance exploration and exploitation, we employ an improved Monte Carlo Tree Search (MCTS) to explore new or underutilized feature space structures after $T$ rounds of iterations. We adopt this MCTS framework because of its random sampling-based search strategy, which enhances the probability of finding superior feature combinations. In our case, this flexible searching method builds the tree on all agents’ generation nodes and adjusts the search direction based on the performance feedback. Here, we define the root node as the original feature set given to the LLM, with each generation node representing a tree node. The search covers all generation nodes in the last iteration layer $Layer_{T}$ as leaf nodes.
Figure 3: The framework of Monte Carlo Tree Search (MCTS) for feature generation includes five stages: 1) Performance Evaluation, 2) Node Selection, 3) Node Expansion, 4) Node Generation, 5) Optimal Subset Searching.
We start with node evaluation by calculating the improvement in the downstream task’s performance compared to the parent node. The evaluation reflects the effectiveness of newly generated features. Then, the MCTS iteratively builds a search tree by cycling through four phases after evaluation: selection, expansion, evaluation, and searching. In the selection phase, children nodes are recursively selected from the root utilizing the Upper Confidence Bound (UCB) (Auer et al., 2002):
$UCB(i)=w_{i}+C\sqrt{\frac{2\ln s^{{}^{\prime}}_{i}}{s_{i}}},$ (5)
where $s_{i}$ and $s^{{}^{\prime}}_{i}$ are the visit counts for node $i$ and its parent respectively, $C$ is a hyperparameter balancing exploration and exploitation. The value $w_{i}$ represents the average performance improvement of all descendant nodes of node $i$, computed based on changes in the downstream task performance metric $\mathbf{\theta}$. Specifically, for each node $i$, the value of $w_{i}$ is calculated as the average effect of performance improvements across all its children in the downstream tasks, given by:
$w_{i}=\frac{1}{n}\sum_{j=1}^{n}(\theta_{t}^{(j)}-\theta_{(t-1)}^{(j)}).$ (6)
Here, $\theta_{t}^{(j)}$ and $\theta_{(t-1)}^{(j)}$ represent the performance metrics at consecutive time points $t$ and $t-1$ for node $j$, indicating the performance of node $j$, and $n$ is the total number of nodes considered.
In the tree, we select the most potential node $s_{l}$. We define the selection criterion for each node as:
$s_{l}^{\text{select}}=\operatorname*{arg\,max}_{s_{l}\in L}\left\{\phi(s_{l})+\psi(s_{l})\right\},$ (7)
where $\psi(s_{l})$ and $\phi(s_{l})$ represent the current and future value of the node. We apply further strategies to generate and expand feature subsets upon selecting these nodes.
On the selected nodes $s_{l}^{\text{select}}$, more complex or novel feature operations are implemented to generate additional feature subsets. These newly generated subsets, $\mathcal{F}_{l}^{{}^{\prime}}$, are then used for further evaluation in downstream tasks to validate their efficacy. The process is mathematically defined as:
$\mathcal{F}_{l}^{\prime}=\mathcal{O}(\mathcal{F}_{l}^{\text{select}}),$ (8)
where $\mathcal{O}$ represents the set of operations applied to transform $\mathcal{F}_{l}^{\text{select}}$ into $\mathcal{F}_{l}^{{}^{\prime}}$. At the end of the MCTS process, the optimal feature subset is determined by choosing the node with the highest value of $w_{i}$ from among the leaf nodes of the tree, representing the feature set that maximized performance improvement during the simulations.
$\mathcal{F}^{*}=\max_{\mathcal{F}_{l}\in\text{Leaf nodes}}w_{l}.$ (9)
This $\mathcal{F}^{*}$, represents the optimal feature subset identified by the MCTS, which should then be validated further in real-world tasks to confirm its efficacy.
4 Experiments
In this section, we present three experiments to demonstrate the effectiveness of the LFG. First, we compare the LFG against several baseline methods on multiple downstream classification tasks. Second, we perform a robustness check on the performance improvement of LFG. Finally, we further study and analyze the iterative performance improvements in the experimental results and discuss the reason.
4.1 Experimental Setup
Datasets. We evaluate the LFG method on four real-world datasets from UCI, including Ionosphere (Ino Sigillito et al. (1989), Amazon Commerce Reviews (Ama) Liu (2011), and Abalone (Aba)  Nash et al. (1995), as well as Diabetes Health Indicators Dataset (Dia) Teboul (2022) from Kaggle. The detailed information is shown in Table 1. For each dataset, we randomly selected $55\%$ of data for training.
Table 1: Datasets Description.
Datasets Samples Features Class
Ion 351 34 2
Ama 1,500 10,000 2
Aba 4177 8 3
Dia 441,455 330 2
Metrics. We evaluate the model performance by the following metrics: Overall Accuracy (Acc) measures the proportion of true results (both true positives and true negatives) in the total dataset. Precision (Prec) reflects the ratio of true positive predictions to all positive predictions for each class. Recall (Rec), also known as sensitivity, reflects the ratio of true positive predictions to all actual positives for each class. F-Measure (F1) is the harmonic mean of precision and recall, providing a single score that balances both metrics.
Downstream Tasks. We apply the LFG model across a range of classification tasks, including Random Forests (RF), Decision Tree (DT), K-Nearest Neighbor (KNN) and Multilayer Perceptrons (MLP). We compare the performance outcomes in these tasks both with and without our method.
Baseline Models. We compare the LFG method with raw data (Raw), the Least Absolute Shrinkage and Selection Operator (Lasso), and Feature Engineering for Predictive Modeling using Reinforcement Learning Khurana et al. (2018) (RL). Here, we set the same operation set for both LFG and RL method to consist of square root, square, cosine, sine, tangent, exp, cube, log, reciprocal, sigmoid, plus, subtract, multiply, and divide. For LLM, we perform all the experiments on OpenAI API, GPT-3.5 Turbo model through the OpenAI platform  [^1] https://openai.com/index/openai-api/.
4.2 Experimental Results
Overall Performance. Table 2 shows the overall results of LFG and the baseline models’ performance. From the table we can see that the LFG method consistently surpasses baseline methods across a variety of metrics and datasets. Specifically, we show the results of LFG in 3 iterations of generation, denoted as LFG-3, to compare with LFG in $T\leq 10$ (LFG). On dataset Ion the LFG achieves the best accuracy of $95.6\%$, which is $6.4\%$ increase over raw data on task RF, and $4.5\%$ over RL method. On dataset Ama, the LFG-3 improves the accuracy of raw data of $59.6\%$ to $61.5\%$ while LFG performs a increase of $4.3\%$ on RF. In precision, the LFG also shows superior performance. For example, on dataset Ion the LFG reaches a precision of $90.6\%$, surpassing the highest precision of $86.7\%$ among all baselines on KNN. This result indicates that the LFG is effective in reducing misclassifications. Regarding the metric recall, the LFG increases the Ion’s recall by $9.8\%$ comparing with vanilla data and by $4.3\%$ comparing with best of baselines on DT task, which demonstrates that LFG brings higher improvement on identifying positive samples. For F1, LFG on Ama improves the data performance from $63.8\%$ to $67.6\%$ in three iterations, and the F1 score increases up tp $68.0\%$. As F1 Score is the harmonic mean of precision and recall, this improvement shows the LFG’s capability to balance between reducing misclassifications and minimizing missed classifications.
Table 2: Overall performance on downstream tasks. The best results are highlighted in bold, and the runner-up results are highlighted in underline. (Higher values indicate better performance.)
Metrics Model RF DT KNN MLP
Acc Ion Ama Aba Dia Ion Ama Aba Dia Ion Ama Aba Dia Ion Ama Aba Dia
Raw 0.892 0.596 0.539 0.859 0.873 0.536 0.489 0.794 0.810 0.548 0.529 0.847 0.892 0.639 0.546 0.862
Lasso 0.892 0.627 0.537 0.860 0.880 0.554 0.491 0.796 0.848 0.533 0.532 0.849 0.905 0.671 0.552 0.863
RL 0.911 0.607 0.559 0.859 0.892 0.548 0.508 0.796 0.829 0.542 0.534 0.848 0.911 0.616 0.558 0.864
LFG-3 0.918 0.615 0.564 0.860 0.905 0.590 0.523 0.800 0.873 0.607 0.541 0.848 0.918 0.676 0.572 0.866
LFG 0.956 0.639 0.564 0.861 0.937 0.590 0.523 0.800 0.886 0.607 0.541 0.848 0.943 0.680 0.573 0.866
Prec Ion Ama Aba Dia Ion Ama Aba Dia Ion Ama Aba Dia Ion Ama Aba Dia
Raw 0.884 0.598 0.533 0.681 0.899 0.537 0.492 0.588 0.867 0.605 0.525 0.643 0.900 0.638 0.541 0.697
Lasso 0.866 0.624 0.531 0.684 0.857 0.554 0.493 0.590 0.865 0.544 0.532 0.647 0.904 0.671 0.549 0.703
RL 0.919 0.610 0.554 0.683 0.892 0.548 0.495 0.590 0.856 0.542 0.528 0.642 0.915 0.616 0.544 0.708
LFG-3 0.915 0.623 0.562 0.688 0.892 0.591 0.524 0.592 0.891 0.608 0.544 0.645 0.940 0.676 0.560 0.719
LFG 0.962 0.639 0.562 0.690 0.929 0.591 0.524 0.592 0.906 0.608 0.544 0.645 0.953 0.681 0.562 0.719
Rec Ion Ama Aba Dia Ion Ama Aba Dia Ion Ama Aba Dia Ion Ama Aba Dia
Raw 0.880 0.597 0.545 0.570 0.837 0.537 0.493 0.597 0.768 0.548 0.533 0.577 0.892 0.639 0.544 0.595
Lasso 0.887 0.622 0.541 0.575 0.874 0.554 0.496 0.599 0.802 0.510 0.542 0.580 0.885 0.672 0.552 0.591
RL 0.911 0.607 0.559 0.568 0.892 0.548 0.508 0.598 0.829 0.542 0.534 0.579 0.911 0.616 0.558 0.589
LFG-3 0.904 0.618 0.564 0.568 0.902 0.591 0.526 0.601 0.828 0.607 0.546 0.582 0.897 0.676 0.573 0.562
LFG 0.944 0.639 0.564 0.570 0.935 0.591 0.526 0.601 0.828 0.607 0.546 0.582 0.921 0.681 0.565 0.562
F1 Ion Ama Aba Dia Ion Ama Aba Dia Ion Ama Aba Dia Ion Ama Aba Dia
Raw 0.882 0.595 0.537 0.587 0.855 0.536 0.492 0.592 0.781 0.476 0.528 0.592 0.889 0.638 0.541 0.618
Lasso 0.875 0.622 0.533 0.594 0.865 0.554 0.493 0.594 0.821 0.403 0.534 0.595 0.893 0.671 0.550 0.614
RL 0.909 0.602 0.556 0.584 0.890 0.548 0.498 0.594 0.814 0.542 0.529 0.594 0.909 0.615 0.544 0.611
LFG-3 0.909 0.612 0.562 0.584 0.897 0.590 0.525 0.596 0.849 0.607 0.541 0.598 0.911 0.676 0.563 0.576
LFG 0.952 0.639 0.562 0.587 0.932 0.590 0.525 0.596 0.854 0.607 0.541 0.598 0.935 0.680 0.563 0.576
Robustness Check. We validate the experiments through five-fold cross-validation. The LFG methods demonstrate robust performance across various datasets and achieve notable metric improvements, including $2.72\%$ in accuracy and approximately $5\%$ in precision, recall and F1 scores with LFG-3, and further $4.81\%$ in accuracy and approximately $7\%$ in other metrics with LFG. The results illustrate the adaptability and effectiveness of the LFG. The extended iterations continue to refine and optimize the feature space, which is reflected in the continuous improvements in precision and recall. The consistent enhancement across all metrics demonstrates the LFG’s reliability and stability comprehensively.
Iterative Performance Improvements. We move forward comparing the performance improvements of LFG-3 and LFG. On one hand, the incremental improvement from LFG-3 to LFG illustrates the continuous enhancement in all metrics, indicating that during additional iterations, the model refines and adjusts its generation strategy to explore better solutions for the feature space. On the other hand, the results of LFG-3 with only three iterations already shows preservative enhancement in all datasets. This shows that the model effectively and rapidly finds the zones of optimal feature space within limited steps of searching. Showing in Figure 5, we can see that the average number of features in a feature subset increases continuously in the first $5$ rounds of generation. The increase in the number of features implies that the the feature generation algorithm is actively finding new, potentially feature combinations in the feature space and is capable of expanding the feature space effectively.
Figure 4: Comparison on KNN (Accuracy and F1).
Figure 5: Increase of Feature Numbers.
5 Conclusion
In this paper we present a novel feature generation method LFG utilizing LLM to enhance the interpretability, adaptability, and strategic flexibility of automated feature engineering. Our target in designing the model is to address the challenges in traditional feature generation including lack of explainability, limited applicability, and rigid strategy formulation with expert-level LLM agents. These problems can significantly limit the broader deployment of feature engineering on machine learning tasks in diverse scenarios. Thus we present LFG with a dynamic and adaptive automated feature generation approach to enhance the interpretability and extends the utility across various data types and tasks. Moreover, we present extensive experimental results to demonstrate that our approach outperforms various existing methods. For future work, we plan to extend this automated feature engineering paradigm across various feature engineering techniques and machine learning tasks, and to conduct a thorough investigation into the transformations that our approach applies to the feature space to study deeper insights into its underlying mechanisms and impacts.
6 Limitations and Ethics Statements
While our LFG shows significant advancements and wide adaptability, there are several limitations that require further exploration, including high computational demands and limited scalability with very large or complex datasets. Second, the effectiveness of the generated features heavily relies on the quality of input data, which can affect the model’s performance in scenarios with poorly curated or noisy datasets. Finally, there may be inherent limitations of adopting LFG in tasks with unique requirements.
The LFG framework can enhance the effectiveness and explainability of the feature generation process with LLM. However, as the framework uses pre-trained GPT-3.5 Turbo as the generation model, it may inherit the ethical concerns associated with GPT-3.5 Turbo, such as responding to harmful queries or exhibiting biased behaviors.
Appendix A Appendix
A.1 Definitions and Symbols
Table 3: Mathematical Symbols in Feature Set and Operation Set.
Symbol Explanation
$\mathcal{F}$ Feature set, represented as =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{n}\}$, where $f_{i}$ is the $i$-th feature.
$\mathcal{D}$ Dataset, represented as a series of samples and their corresponding labels, $\mathcal{D}=\langle\mathcal{F},\mathcal{y}\rangle$.
$\mathcal{y}$ Target label set, corresponding to the samples in the dataset $\mathcal{D}$.
$\mathcal{O}$ Operation set, containing all available mathematical operations, $\mathcal{O}=\mathcal{O}_{\text{unary}}\cup\mathcal{O}_{\text{binary}}$.
$\mathcal{O}_{\text{unary}}$ Unary operation set, e.g., $\mathcal{O}_{\text{unary}}=\{f_{1}^{2},e^{f_{1}},\log(f_{1}),\ldots\}$.
$\mathcal{O}_{\text{binary}}$ Binary operation set, e.g., $\mathcal{O}_{\text{binary}}=\{f_{1}+f_{2},f_{1}\times f_{2},f_{1}/f_{2},\ldots\}$.
$\mathcal{R}$ Downstream machine learning task (e.g., classification, regression, ranking, detection).
$\phi$ Performance indicator of the downstream task $\mathcal{R}$.
$\hat{\mathcal{F}}$ Updated feature set, derived by applying operations from $\mathcal{O}$ to the original feature set $\mathcal{F}$.
A.2 Definitions of Feature Operations
[…]
    $\text{square}(f_{i})=f_{i}^{2}$ (10)
[…]
    $\text{exp}(f_{i})=e^{f_{i}}$ (11)
[…]
    $\text{log}(f_{i})=\log(f_{i})$ (12)
[…]
    $\text{sin}(f_{i})=\sin(f_{i})$ (13)
[…]
    $\text{cos}(f_{i})=\cos(f_{i})$ (14)
[…]
    $\text{plus}(f_{i},f_{j})=f_{i}+f_{j}$ (15)
[…]
    $\text{multiply}(f_{i},f_{j})=f_{i}\times f_{j}$ (16)
[…]
    divide$(f_{i},f_{j})=\frac{f_{i}}{f_{j}}$ (17)
Figure 6: The exemplary prompt for LLM to generate agents and features. The Feature Engineering is shown in blue, the Iterative Refinement and Evaluation is shown in green and the Markdown and Organization is shown in purple. The LLM follows the instructions to start the first iteration and continues to generate features with downstream tasks’ performance data.
A.3 Prompting Structure
At the beginning of the iteration, we provide the LLM with a prompt to guide it in generating features as expected. The prompt consists of three parts: (1) Feature Engineering, (2) Iterative Refinement and Evaluation, and (3) Markdown and Organization. Here are the detailed descriptions of these parts:
  1. 1.
    Feature Engineering. In the first round of iteration, $l$ agents independently employ the Tree of Thoughts method on the original feature set $\mathcal{F}$. They select, apply operations to, or delete features to generate new features and form new feature subsets $\mathcal{F}_{l}$.
  2. 2.
    Iterative Refinement and Evaluation. Then, we submit these $l$ feature subsets to downstream tasks to assess the performance of each $\mathcal{F}_{l}$, defined by the indicators $\mathbf{\theta_{1}}=(\theta^{1},\ldots,\theta^{l})$. We provide $\mathbf{\theta_{1}}$ and the performance of the original feature set $\mathbf{\theta_{0}}$ in the downstream tasks as feedback. Each expert shares their thought process in detail, including considering the ideas of others and acknowledging errors, and evaluates their performance based on the improvement in the downstream tasks.
  3. 3.
    Markdown and Organization. We organize the entire process in a Markdown table format, meticulously documenting every step of selection, operation, and deletion. As we enter the second iteration, each agent, based on their respective feature subset $\mathcal{F}_{l}$, generates new features guided by the evaluation and downstream task metrics.
In the prompt, we implement the Tree of Thoughts (ToT) methodology. Here, the process that each agent analyzes and selects features to operate on plays the role of Thought Decomposition. The key to this generation process is logical analysis and the prediction of potential impacts of features of the whole feature set. In this process, agents generate multiple potential solutions according to the current state of the feature set. At the end of each round, we evaluate the performance of new feature subsets in downstream tasks. This evaluation and the comparison of performance before and after play the role of the State Evaluator. This process provides a quantitative evaluation for the iterative optimization of the feature set. Through several iterations, the entire LLM continuously adjusts and optimizes, thus searching for the optimal feature set that can significantly enhance the performance of downstream tasks.